I'm really not sure what topic this comes under but how do you calculate these sums?


Given that 76 × 219 = 16644 work out the value of
(a) 16644 ÷ 21.9
(b) 75 × 219
Thanks.

Answer by Alan3354(69443)

Given that 76 × 219 = 16644
That's true, tho not relevant.
2*2 = 4, also.
----------
work out the value of
(a) 16644 ÷ 21.9
Hint: 16644/219 = 76
=============================
(b) 75 × 219
= 76*219 - 1*219
--------------
PS Neither or those is a sum.